Peroni Company paid wages of $167,700 this year. Of this amount, $101,100 was taxable for net FUTA and SUTA purposes. The state's contribution tax rate is 3.1% for Peroni Company. Due to cash flow problems, the company did not make any SUTA payments until after the Form 940 filing date. Compute the following; round your answers to the nearest cent.

a. Amount of credit the company would receive against the FUTA tax for its SUTA contributions
$

b. Amount that Peroni Company would pay to the federal government for its FUTA tax
$

c. Amount that the company lost because of its late payments
$
Answer #1

Important points to note:

1) The FUTA tax rate applies to the first \$ 7,000 of wages paid to each employee, the rate being 6%.

2) Credit against FUTA tax is allowable upto 5.4%.

3) If the SUTA dues are paid on time, SUTA contributions deductions will be allowed.

4) In case of any delay, credit equalling to 90% of SUTA dues is allowable.

5) The due date for filing Form 940 is on 31st January. However, if all FUTA dues are deposited, there is an extension upto February 10th.

6) In case of any delay in payments, the following charges will be incurred -

(A) Penalty of 10% of quarterly premium amount

(B) 1% interest on the premium amount due for each month which remains outstanding

(C) And, in case of deliberate fraud, a penalty upto 50% of outstanding payment.

Keeping this in mind, we can move forward:

a. Given, State's contribution tax rate( i.e., SUTA) = 3.1% & Taxable wages = \$ 101,100

\therefore SUTA contributions = 3.1% of \$ 101,100 = \$ 3,134 (rounded)

However, since there was a delay in SUTA payments, only 90% of SUTA contributions would be allowed as credit against FUTA tax = 90% of \$ 3,134 = \$ 2,821 (rounded).

----------------------------------------------------------------------------------------------------------------------------------------------------

b. Total taxable wages (i) = \$ 101,100

FUTA tax rate (ii) = 6%

Amount of Gross FUTA Tax [(iii) = (i) * (ii)] = \$ 6,066

Credit against FUTA tax (iv) = 5.4%

Amount of credit [(v) = (i) * (iv)] = \$ 5,459 (rounded)

\therefore Amount of FUTA tax [(vi) = (iii) - (v)] = \$ 607

----------------------------------------------------------------------------------------------------------------------------------------------------

c. Since Peroni company makes SUTA payments after January 31st, the company incurs 10% penalty alongwith 1% interest of SUTA dues [Refer Point 6 (A) & 6(B)]

= (10% of \$ 3,134) + (1% of \$ 3,134)

= \$ 313 + \$ 31 = \$ 344

Also, 10% of SUTA dues will not be allowed as credit due to delay, i.e., \$ 313.

\therefore Total loss to company = \$ 344 + \$ 313 = \$ 657.

----------------------------------------------------------------------------------------------------------------------------------------------------